brent lent his brother bob some money at 8% simple interest and he lent his sister betty half as much money at twice the interest rate. both loans were for one year. if brent made a total of 24 cents in interest then how much did he lend to each one

Answer by stanbon(75887)

brent lent his brother bob some money at 8% simple interest and he lent his sister betty half as much money at twice the interest rate. both loans were for one year. if brent made a total of 24 cents in interest then how much did he lend to each one
----
Interest + Interest = Interest
0.08x + 0.16(x/2) = 0.24 dollars
-----
8x + 8x = 24
16x = 24
x = 3/2 = $1.50 (amt. leant to bob)
x/2 = 3/4 = $0.75 (amt. leant to sister)
------
Check::
0.08*1.5 + 0.16(0.75) = 0.24 dollars = 24 cents
-------------------
Chgeers,
Stan H.
----------------